Oprah Winfrey has closed on a 42-acre estate near Santa Barbara, California, for $49,800,000. If Oprah puts 30% down and finances at 7% for 30 years, what would her monthly payment be? (Do not round intermediate calculations. Round your answer to the nearest dollar amount.)
Answer #1
Monthly Payment = Loan amount / Present value of annuity of 1
= $ 3,48,60,000 / 150.30757
= $       2,31,924
Working:
# 1
Loan amount = Cost of automobile * (1- Down Payment percentage)
= $ 4,98,00,000 * (1-0.30)
= $ 3,48,60,000
# 2
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.005833)^-360)/0.005833 i = 7%/12 = 0.005833
= 150.3075679 n = 30*12 = 360